3.You borrow $100,000 today, the annual interest rate is 12%, you repay the interest and principle once a year. The loan will be payoff in 30 years. What is your annual payment? 4.Matt plans to make 35 equal yearly deposits into his retirement account starting one vear from now (t-1). Starting at t-36, he will make 20 equal withdrawals of S150,000 each year from this account. The intere must Matt deposit each year to satisfy his retirement needs? 5.Kevin is planning on saving for retirement. He wishes to have $500,000 in his retirement account in 40 years (t-40). He plans on doing this by making equal yearly deposits into his savings account starting from t-1. The interest rate is 4% pa. How much does he need to deposit each year? (20 points) st rate is 7% pa. Ho w much
Answer #1
As per HOMEWORKLIB POLICY only 1 question per submission can be solved. Here comes the solution for first question
PV of annuity for making pthly payment
P = PMT x (((1-(1 + r) ^- n)) / i)
Where:
P = the present value of an annuity stream
PMT = the dollar amount of each annuity payment
r = the effective interest rate (also known as the discount rate)
i=nominal Interest rate
n = the number of periods in which payments will be made
PV 100000
Time 30
Fixed rate 12%
100000 = PMT x (((1-(1 + r) ^- n)) / i)
100000 = PMT * (((1-(1 + 12%) ^- 30)) / 12%)
100000 = PMT * 8.05518
PMT= 100000/8.05518
PMT=     12,414.37